OpenStudy (anonymous):

Given the equation 5x+10= -30, solve for x and choose the correct solution. x=-8 x=8 x=-4 x=4

OpenStudy (anonymous):

5x+10=-30 -10 -10 5x=-40 /5 /5 x= -8

OpenStudy (anonymous):

Ty.